# VoiceShop: A Unified Speech-to-Speech Framework for Identity-Preserving Zero-Shot Voice Editing

### 10 Apr 2024 v1

@@ Abstract
We present VoiceShop, a novel speech-to-speech framework that can modify multiple attributes of speech, such as age, gender, accent, and speech style, in a single forward pass while preserving the input speaker’s timbre. Previous works have been constrained to specialized models that can only edit these attributes individually and suffer from the following pitfalls: the magnitude of the conversion effect is weak, there is no zero-shot capability for out-of-distribution speakers, or the synthesized outputs exhibit timbre leakage which changes the speaker’s perceived identity. Our work proposes solutions for each of these issues in a simple modular framework based on a conditional diffusion backbone model with optional normalizing flow-based and sequence-to-sequence speaker attribute-editing modules, whose components can be combined or removed during inference to meet a wide array of tasks without additional model finetuning. Audio samples are available at https://voiceshopai.github.io.
Audio Synthesis, Speech Synthesis, Speech-to-Speech, Zero-Shot, Disentangled Representation Learning, Diffusion Model, Flow-Based Generative Model, Seq2Seq, BN2BN, Voice Conversion, Accent Conversion, Speech Style Conversion, Age Conversion, Gender Conversion
@@ 1 Introduction
Even as the quality of proposed speech synthesis models improves following advances in score-based diffusion (Kong et al., 2020b) and neural codec modeling (Défossez et al., 2022; Zeghidour et al., 2021), disentangled representation learning, whereby attributes such as a speaker’s timbre, prosody, age, gender, and accent are extracted and separated from speech signals, remains an open problem (Peyser et al., 2022; Polyak et al., 2021; Wang et al., 2023c). Disentangling these characteristics is an essential component of any generative model that aims to allow users to flexibly modify certain attributes of their speech while keeping others constant, all while preserving their identity, i.e., their timbre. Due to the non-trivial nature of this task, most works focus on editing a single attribute at a time and frequently require that synthesized speech map to a set of in-domain target speakers seen during training, leaving the development of speech systems capable of editing multiple attributes at once under-explored. Therefore, the prevailing limitation of existing models is that such fine-grained control of generated speech in a user’s own voice is not provided.
To this end, we present VoiceShop, a novel speech-to-speech foundation model capable of a wide assortment of zero-shot synthesis tasks. Within a unified framework, VoiceShop is capable of monolingual and cross-lingual voice conversion, identity-preserving many-to-many accent and speech style conversion, and age and gender editing, where all tasks may be performed in a zero-shot setting on arbitrary out-of-domain speakers not seen during training. Furthermore, our framework enables users to perform any combination of these tasks simultaneously in a single forward pass while preserving the original speaker’s identity, i.e., one may edit the age, gender, accent, and speech style of input speech at once. To our knowledge, VoiceShop is among the first proposed methods to be capable of performing combined multi-attribute editing across languages in this manner.
@@
This behavior is achieved through the use of a diffusion backbone model, which accepts global speaker embeddings and time-varying content features as conditioning signals, enabling robust zero-shot VC. We additionally train two separate task-specific editing modules based on a continuous normalizing flow model (Rezende & Mohamed, 2016) that operates on the global speaker embedding to achieve age and gender editing and a sequence-to-sequence model (Sutskever et al., 2014) that operates on the local content features to achieve accent and speech style conversion. Together, these components make up VoiceShop’s modular framework and enable its flexible VE capabilities without finetuning. Our main contributions are as follows:
  • Multi-attribute speech editing via unified scalable framework: Our speech-to-speech framework supports both conventional VC and VE in a single forward pass. Editing modules are trained separately from the backbone synthesis module and may be used in a modular plug-and-play fashion, making them scalable to novel editing tasks and datasets without finetuning.
  • Zero-shot generalization: Our model achieves strong zero-shot VC and VE performance on unseen speakers, matching or outperforming existing state-of-the-art (SOTA) baselines specialized for different sub-tasks.
@@ 2 Related Work
Figure 1: The architecture of VoiceShop: The overall pipeline follows an analysis-synthesis approach. During the analysis step, the ECAPA-TDNN speaker encoder and pre-trained ASR module respectively decompose input speech into speaker identity represented by a global speaker embedding (SE) and local content embeddings represented by a sequence of bottleneck (BN) features. During the synthesis step, both SE and BN features condition the diffusion backbone model to reconstruct mel-spectrograms of input speech, followed by a vocoder to acquire time-domain waveforms. Voice editing: Editing is an optional step in between the analysis and synthesis steps reserved for inference. As highlighted in the two dashed boxes, an attribute-conditional flow module is used to globally edit the speaker embedding (e.g., change age and gender), whereas a BN2BN module is used to edit content embeddings (e.g., change the prosody or speech style). These voice editing modules are trained separately from the generative backbone module and are used in a modular plug-and-play manner.
@@
Zero-Shot Voice Conversion. While fully supervised VC is considered a canonical problem in the speech domain, AutoVC (Qian et al., 2019) is among the earliest works to successfully extend VC to the zero-shot setting, while demonstrating promising results could be achieved using a simple autoencoder architecture and carefully designed information bottleneck. Since this work, a variety of methods has been proposed to improve the generalization abilities of zero-shot any-to-any VC systems, such as end-to-end normalizing flow models that maximize the variational lower bound of the marginal log-likelihood of data (Casanova et al., 2022; Kim et al., 2021; Li et al., 2022a; Pankov et al., 2023), unsupervised mutual information-based strategies for speaker-content disentanglement (Wang et al., 2021), and diffusion-based approaches (Choi et al., 2023; Kim et al., 2023; Popov et al., 2021). While many of these proposals achieve high-quality outputs, several do not acknowledge the cross-lingual case (Casanova et al., 2022), whereby source and target utterances are spoken in different languages. In this work, we show that it is possible to achieve strong zero-shot VC performance using a relatively simple diffusion modeling approach while supporting both monolingual and cross-lingual conversion.
@@
Accent and Speech Style Conversion. Style transfer refers to the task of disentangling stylistic and semantic information of samples from a given domain in order to synthesize new samples which impose the style of a reference sample, while retaining the content of another (Gatys et al., 2015). Accent conversion (AC) or speech style conversion refines this definition to focus on converting the accent or “speech style” of an utterance, i.e., modifying a speaker’s pronunciation or prosody, while preserving their spoken content and timbre. Several works have been proposed to achieve accent or speech style conversion, but are often limited by any combination of the following factors. The models may lack zero-shot capabilities, requiring that input timbres are mapped to a set of in-domain target timbres seen during training, altering the identity of the source speaker (Zhang et al., 2022b, a), or rely on adversarial training strategies, such as domain adaptation via gradient reversal introduced by Ganin et al. (2016), or empirically designed data augmentation methods to encourage disentanglement (Badlani et al., 2023; Chan et al., 2022; Choi et al., 2022; Jia et al., 2023; Li et al., 2022b, c; Zhang et al., 2023a). Many frameworks are designed for text-to-speech (TTS), but do not address the speech-to-speech (STS) case (Guan et al., 2023; Tinchev et al., 2023; Zhou et al., 2023), require ground truth alignments between input and output features acquired through third-party forced alignment tools (Ezzerg et al., 2022; Karlapati et al., 2022), lack explicit control of specific attributes (Wang et al., 2018), or do not support many-to-many conversion, i.e., one must train separate models for each desired target accent, rather than encompassing all conversion paths in a single model (Zhao et al., 2018, 2019). To our knowledge, no works have been proposed for cross-lingual AC, whereby accents extracted from speech of one language are transferred to speech of another language.
@@
Age and Gender Editing. Within the context of VE, NANSY++ (Choi et al., 2022) is the most relevant work to support age and gender editing in a zero-shot manner, achieved through a unified framework for synthesizing and manipulating speech signals from analysis features. In their voice designing pipeline NANSY-VOD, the authors employ three normalizing flow networks in a cascaded manner, where the outputs of the earlier models are passed as conditions to subsequent models, to predict $F_{0}$ statistics, a global timbre embedding, and a fixed number of timbre tokens used to edit the age and gender of output speech. We show that VoiceShop achieves age and gender editing capabilities using a simpler design consisting of one normalizing flow model that allows simultaneous attribute editing.
@@ 3.1 Method Overview
The first two stages listed above consist of large-scale pre-training of the diffusion backbone model to achieve robust zero-shot VC ability. The third stage focuses on separately preparing lightweight modules that respectively operate on the diffusion model’s two conditioning signals in order to modify one or more speech attributes such as age, gender, accent, or speaking style during inference. By tackling VE in this modular approach, we remove the need for transfer learning via finetuning on downstream tasks later on.
@@ 3.2.2 Conditional Diffusion Backbone Model for Zero-Shot Voice Conversion
Table 1: Datasets used to train our conditional diffusion backbone model. Monolingual models are only trained on the English subset, whereas bilingual models are trained on all listed data.
@@
We propose a conditional diffusion model to predict mel-spectrogram representations of speech, serving as the backbone of our unified framework. Specifically, we consider a forward process $p_{\theta}(x_{0}|E_{S},E_{C})$, where $x_{0}\in\mathbb{R}^{F\times L}$ denotes a mel-spectrogram extracted from raw audio, such that $F$ is the number of frequency bands and $L$ refers to duration, $S(x_{0})=E_{S}\in\mathbb{R}^{D_{S}}$ denotes an utterance-level global speaker embedding (i.e., lacking temporal information) produced by speaker encoder $S$, and $C(x_{0})=E_{C}\in\mathbb{R}^{D_{C}\times\frac{L}{4}}$ denotes a time-varying content embedding produced by a pre-trained ASR or APR model $C$. We train the speaker encoder jointly with the diffusion model from scratch, adopting the same configuration as the ECAPA-TDNN speaker verification model (Desplanques et al., 2020) to extract timbre information as a 512-dimensional vector.
As with Moûsai (Schneider et al., 2023), we employ a one-dimensional U-Net (Ronneberger et al., 2015) as the architecture of the diffusion model and modify their open-source implementation based on the A-UNet toolkit [^2] https://github.com/archinetai/audio-diffusion-pytorch. We follow their design for each U-Net block with some modifications, the details of which are discussed in §A.2. We also adopt the velocity-based formulation of the diffusion training objective and angular reparametrization of DDIM (Song et al., 2022) proposed by Salimans & Ho (2022), defining our forward process as:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\alpha_{t}x_{0},\beta_{t}\epsilon)$ (1)
$\mathcal{L}(\theta)$ $\coloneqq\mathbb{E}_{t\sim\mathcal{U}[0,1]}\left[\|v_{t}-\hat{v}_{t}(x_{t};t,E_{S},E_{C})\|_{2}^{2}\right]$ (2)
$v_{t}$ $=\alpha_{t}\epsilon-\beta_{t}x_{0}$ (3)
$\hat{x}_{0}$ $=\alpha_{t}x_{t}-\beta_{t}\hat{v}_{t}$ (4)
$\hat{\epsilon}_{t}$ $=\beta_{t}x_{t}+\alpha_{t}\hat{v}_{t}$ (5)
$\hat{x}_{t-1}$ $=\alpha_{t-1}\hat{x}_{0}+\beta_{t-1}\hat{\epsilon}_{t}$ (6)
[…]
Table 2: Training configurations of each diffusion backbone model.
Backbone Model ASR Model Layer Bilingual
VS-EN-L10 ASR-EN 10
VS-EN-L18 ASR-EN 18
VS-EN-CN-L10 ASR-EN-CN 10
VS-EN-CN-L18 ASR-EN-CN 18
We train all diffusion backbone models using the AdamW optimizer (Loshchilov & Hutter, 2019) on 4 A100 GPUs with a learning rate of $1\times 10^{-4}$ and batch size of 88 samples for 250K iterations. We set a learning rate decay of 0.85 applied every 20 epochs. To leverage large amounts of speech data, we train with unlabeled audio in a self-supervised manner. After convergence, the diffusion backbone model becomes capable of robust zero-shot VC.
@@ 3.3.1 Attribute-Conditional Normalizing Flow for Age and Gender Editing
We observe that the speaker encoder jointly trained with our diffusion backbone model achieves strong speaker identity disentanglement, indicating that many attributes like age and gender are encoded into the global 512-dimensional speaker embedding vector. Therefore, the manipulation of these attributes can be seen as re-sampling from the learned latent space of speaker embeddings. To achieve fully controllable generation and editing of specific speaker attributes while leaving other attributes unaffected, we take inspiration from StyleFlow (Abdal et al., 2021) in the image editing domain and employ a similar attribute-conditional continuous normalizing flow (CNF) that operates on this latent space.
@@
The training objective of the CNF would be to maximize the likelihood of the data $\mathbf{w}$. In our attribute conditional scenario, given the speaker attributes vector $\mathbf{a}$ (e.g., for age and gender, $\mathbf{a}\in\mathbb{R}^{2}$) that is associated with sample $\mathbf{w}$, we update the ODE network to be $\phi_{\theta}(z(t),\mathbf{a},t)$, which is conditioned on both $t$ and $\mathbf{a}$. Consequently, our new objective becomes $\max_{\theta}\sum_{\mathbf{w},\mathbf{a}}\log p(\mathbf{w}|\mathbf{a},t)$. During training, we also employ trajectory polynomial regularization (Huang & Yeh, 2021), which we find stabilizes the training. Note that there is no attribute editing during the training process. A well-trained CNF model behaves as a bijective mapping between the data distribution and prior distribution with the help of an ODE solver. In our experiments, we use the dopri5 ODE solver (Hairer et al., 2008) and adopt the same CNF implementation from StyleFlow [^3] https://github.com/RameenAbdal/StyleFlow. We only use one CNF block with hidden dimension 512 rather than stacking multiple blocks, which leads to lower validation loss. Our lightweight CNF model has 0.5M learnable parameters and is trained using the Adam optimizer (Kingma & Ba, 2014) on a single V100 GPU for 24H. The attribute editing procedure at inference time is depicted in Figure 2.
Attributes Dataset. The attribute labels for training our attribute-conditional CNF are obtained from the CommonVoice 13.0 dataset (Ardila et al., 2019), which contains multiple audio recordings of more than 51K human-validated anonymous speakers, totaling 2,429 hours. Among these validated speakers, about 20K speakers provide age (from twenties to nineties) and gender labels (male or female). To fully utilize the dataset, we first train a naive ECAPA-TDNN-based age and gender prediction model by appending a projection layer that predicts age as a numeric value and gender as a logit to its original output layer. We combine mean absolute error (MAE) and cross entropy (CE) losses for age and gender prediction respectively. This prediction model serves two purposes:

### 11 Apr 2024 v2

@@ Abstract
We present VoiceShop, a novel speech-to-speech framework that can modify multiple attributes of speech, such as age, gender, accent, and speech style, in a single forward pass while preserving the input speaker’s timbre. Previous works have been constrained to specialized models that can only edit these attributes individually and suffer from the following pitfalls: the magnitude of the conversion effect is weak, there is no zero-shot capability for out-of-distribution speakers, or the synthesized outputs exhibit undesirable timbre leakage. Our work proposes solutions for each of these issues in a simple modular framework based on a conditional diffusion backbone model with optional normalizing flow-based and sequence-to-sequence speaker attribute-editing modules, whose components can be combined or removed during inference to meet a wide array of tasks without additional model finetuning. Audio samples are available at https://voiceshopai.github.io.
**footnotetext: Equal contribution
@@ 1 Introduction
Even as the quality of proposed speech synthesis models improves following advances in diffusion (Kong et al., 2020b) and neural codec modeling (Défossez et al., 2022; Zeghidour et al., 2021), disentangled representation learning, whereby attributes such as a speaker’s timbre, prosody, age, gender, and accent are extracted and separated from speech signals, remains an open problem (Peyser et al., 2022; Polyak et al., 2021; Wang et al., 2023c). Disentangling these characteristics is an essential component of any generative model that aims to allow users to flexibly modify certain attributes of their speech while keeping others constant. Due to the non-trivial nature of this task, several works focus on editing a single attribute at a time and frequently require that synthesized speech map to a set of in-domain target speakers seen during training, leaving the development of systems capable of editing multiple attributes at once under-explored. Therefore, the prevailing limitation of existing models is that such fine-grained control of generated speech in a user’s own voice is limited.
To this end, we present VoiceShop, a novel speech-to-speech foundation model capable of a wide assortment of zero-shot synthesis tasks. Within a unified framework, VoiceShop is capable of monolingual and cross-lingual voice conversion, identity-preserving many-to-many accent and speech style conversion, and age and gender editing, where all tasks may be performed in a zero-shot setting on arbitrary out-of-domain speakers not seen during training. Furthermore, our framework enables users to perform any combination of these tasks simultaneously in a single forward pass while preserving the original speaker’s unedited attributes, i.e., one may edit the age, gender, accent, and speech style of input speech at once.
@@
This behavior is achieved through the use of a diffusion backbone model, which accepts global speaker embeddings and time-varying content features as conditioning signals, enabling robust zero-shot VC. We additionally train two separate task-specific editing modules based on a continuous normalizing flow model (Rezende and Mohamed, 2016) that operates on the global speaker embedding to achieve age and gender editing and a sequence-to-sequence model (Sutskever et al., 2014) that operates on the local content features to achieve accent and speech style conversion. Together, these components make up VoiceShop’s modular framework and enable its flexible VE capabilities without finetuning. Our main contributions are as follows:
  • Multi-attribute speech editing via unified scalable framework: VoiceShop supports both conventional VC and VE in a single speech-to-speech framework. Editing modules are trained separately from the backbone synthesis module and may be used in a modular plug-and-play fashion, making them scalable to novel editing tasks and datasets without additional finetuning.
  • Zero-shot generalization: Our model achieves strong zero-shot VC and VE performance on arbitrary unseen speakers, matching or outperforming existing state-of-the-art (SOTA) baselines specialized for different sub-tasks.
  • Disentangled attribute control: We achieve strong disentanglement of multiple speaker attributes, which in the VE case allows simultaneous identity-preserving transformations of users’ voices without modifying their timbre.
@@ 2 Related Work
Figure 1: The architecture of VoiceShop: The overall pipeline follows an analysis-synthesis approach. During the analysis step, the ECAPA-TDNN speaker encoder and pre-trained ASR module respectively decompose input speech into speaker identity represented by a global speaker embedding (SE) and local content embeddings represented by a sequence of bottleneck (BN) features. During the synthesis step, both SE and BN features condition the diffusion backbone model to reconstruct mel-spectrograms of input speech, followed by a vocoder to acquire time-domain waveforms. Voice editing: Editing is an optional step in between the analysis and synthesis steps reserved for inference. As highlighted in the two dashed boxes, an attribute-conditional flow module is used to globally edit the speaker embedding (e.g., change age and gender), whereas a BN2BN module is used to edit content embeddings (e.g., change prosody or speech style). These voice editing modules are trained separately from the generative backbone module and are used in a modular plug-and-play manner.
Zero-Shot Voice Conversion.
@@
While fully supervised VC is considered a canonical problem in the speech domain, AutoVC (Qian et al., 2019) is among the earliest works to successfully extend VC to the zero-shot setting, demonstrating that promising results could be achieved using a simple autoencoder architecture and carefully designed information bottleneck. Since this work, a variety of methods has been proposed to improve the generalization abilities of zero-shot any-to-any VC systems, such as end-to-end normalizing flow models that maximize the variational lower bound of the marginal log-likelihood of data (Casanova et al., 2022; Kim et al., 2021; Li et al., 2022a; Pankov et al., 2023), unsupervised mutual information-based strategies for speaker-content disentanglement (Wang et al., 2021), and diffusion-based approaches (Choi et al., 2023; Kim et al., 2023; Popov et al., 2021). While many of these proposals achieve high-quality outputs, the cross-lingual case (Casanova et al., 2022), whereby source and target utterances are spoken in different languages, is less explored. In this work, we show that it is possible to achieve strong zero-shot VC performance using a relatively simple diffusion modeling approach while supporting monolingual and cross-lingual conversion.
Accent and Speech Style Conversion.
@@
Style transfer refers to the task of disentangling stylistic and semantic information of samples from a given domain in order to synthesize new samples that impose the style of a reference sample, while retaining the content of another (Gatys et al., 2015). Accent conversion (AC) or speech style conversion refines this definition to focus on converting the accent or “speech style” of an utterance, i.e., modifying a speaker’s pronunciation or prosody, while preserving their spoken content and timbre. Several works have been proposed to achieve accent or speech style conversion, but are often limited by any combination of the following factors. The models may lack zero-shot capabilities, requiring that input timbres are mapped to a set of in-domain target timbres seen during training, altering the identity of the source speaker (Zhang et al., 2022b, a), or rely on adversarial training strategies, such as domain adaptation via gradient reversal introduced by Ganin et al. (2016), or empirically designed data augmentation methods to encourage disentanglement (Badlani et al., 2023; Chan et al., 2022; Choi et al., 2022; Jia et al., 2023; Li et al., 2022b, c; Zhang et al., 2023a). Many frameworks are designed for text-to-speech (TTS), but do not address the speech-to-speech (STS) case (Guan et al., 2023; Tinchev et al., 2023; Zhou et al., 2023), require ground truth alignments between input and output features acquired through third-party forced alignment tools (Ezzerg et al., 2022; Karlapati et al., 2022), lack explicit control of specific attributes (Wang et al., 2018), or do not support many-to-many conversion, i.e., one must train separate models for each desired target accent, rather than encompassing all conversion paths in a single model (Zhao et al., 2018, 2019). To our knowledge, no works have been proposed for cross-lingual AC, whereby accents extracted from speech of one language are transferred to speech of another language.
Age and Gender Editing.
@@
Within the context of VE, NANSY++ (Choi et al., 2022) is the most relevant work to support age and gender editing in a zero-shot manner, achieved through a unified framework for synthesizing and manipulating speech signals from analysis features. In their voice designing pipeline NANSY-VOD, the authors employ three normalizing flow networks in a cascaded manner, where the outputs of the earlier models are passed as conditions to subsequent models, to predict $F_{0}$ statistics, a global timbre embedding, and a fixed number of timbre tokens used to edit the age and gender of output speech. We show that VoiceShop achieves age and gender editing capabilities using a simpler design consisting of one normalizing flow model that allows simultaneous attribute editing.
@@ 3.1 Method Overview
The first two stages listed above consist of large-scale pre-training of the diffusion backbone model to achieve robust zero-shot VC ability. The third stage focuses on separately preparing lightweight modules that respectively operate on the diffusion model’s two conditioning signals to modify one or more speech attributes such as age, gender, accent, or speaking style during inference. By tackling VE in this modular approach, we remove the need for transfer learning via finetuning on downstream tasks later on.
@@ 3.2.2 Conditional Diffusion Backbone Model for Zero-Shot Voice Conversion
We propose a conditional diffusion model to predict mel-spectrogram representations of speech, serving as the backbone of our unified framework. Specifically, we consider a reverse process $p_{\theta}(x_{0:T}|E_{S},E_{C})$, where $x_{0}\in\mathbb{R}^{F\times L}$ denotes a mel-spectrogram extracted from raw audio, such that $F$ is the number of frequency bands and $L$ refers to the duration, $S(x_{0})=E_{S}\in\mathbb{R}^{D_{S}}$ denotes an utterance-level global speaker embedding (i.e., lacking temporal information) produced by speaker encoder $S$, and $C(x_{0})=E_{C}\in\mathbb{R}^{D_{C}\times\frac{L}{4}}$ denotes a time-varying content embedding produced by a pre-trained ASR or APR model $C$. We train the speaker encoder jointly with the diffusion model from scratch, adopting the same configuration as the ECAPA-TDNN speaker verification model (Desplanques et al., 2020) to extract timbre information as a 512-dimensional vector.
As with Moûsai (Schneider et al., 2023), we employ a one-dimensional U-Net (Ronneberger et al., 2015) as the architecture of the diffusion model and modify their open-source implementation based on the A-UNet toolkit [^2] https://github.com/archinetai/audio-diffusion-pytorch. We follow their design for each U-Net block with some modifications, the details of which are discussed in §A.2. We also adopt the velocity-based formulation of the diffusion training objective and angular reparametrization of DDIM (Song et al., 2022) proposed by Salimans and Ho (2022), defining our forward process as:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\alpha_{t}x_{0},\beta_{t}\epsilon)$ (1)
$\mathcal{L}(\theta)\coloneqq{}$ $\mathbb{E}_{t\sim\mathcal{U}[0,1]}\left[\|v_{t}-f_{\theta}(x_{t};t,E_{S},E_{C})\|_{2}^{2}\right]$ (2)
$v_{t}={}$ $\alpha_{t}\epsilon-\beta_{t}x_{0}$ (3)
$\hat{v}_{t}$ $=f_{\theta}(x_{t};t,E_{S},E_{C})$ (4)
$\hat{x}_{0}$ $=\alpha_{t}x_{t}-\beta_{t}\hat{v}_{t}$ (5)
$\hat{\epsilon}_{t}$ $=\beta_{t}x_{t}+\alpha_{t}\hat{v}_{t}$ (6)
$\hat{x}_{t-1}$ $=\alpha_{t-1}\hat{x}_{0}+\beta_{t-1}\hat{\epsilon}_{t}$ (7)
When deciding which content features to use for conditioning the diffusion backbone model, we consider observations made by Yang et al. (2023b), who find that the selection of which layer to extract content representations from pre-trained ASR or APR models has a measurable impact on the magnitude of various information sources encoded in the latent sequence. Our work validates this finding, and we note that the activation maps of shallower layers of such models, such as ${}^{\text{th}}$ layer, contain far greater amounts of prosodic and accent information compared to those of deeper layers, such as ${}^{\text{th}}$ layer. Intuitively, this phenomenon may be explained by a layer’s proximity to the final loss calculation, such that deeper layers are in some sense “closer” to pure textual transcriptions of input speech, while shallower layers still contain significant amounts of timbre and prosody leakage desirable for accent and speech style conversion.
[…]
Table 1: Datasets used to train diffusion backbone models. Monolingual models only use English data, whereas bilingual models use all listed data.
Table 2: Training configurations of each diffusion backbone model. Layer numbers (10 and 18) are contained in model names.
@@
Backbone Model ASR Model Bilingual
VS-EN-L10 ASR-EN
VS-EN-L18 ASR-EN
VS-EN-CN-L10 ASR-EN-CN
VS-EN-CN-L18 ASR-EN-CN
Table 2: Training configurations of each diffusion backbone model. Layer numbers (10 and 18) are contained in model names.
We train all diffusion backbone models using the AdamW optimizer (Loshchilov and Hutter, 2019) on 4 A100 GPUs with a learning rate of $1\times 10^{-4}$ and batch size of 88 samples for 250K iterations. We set a learning rate decay of 0.85 applied every 20 epochs. To leverage large amounts of speech data, we train with unlabeled audio in a self-supervised manner. After convergence, the diffusion backbone model becomes capable of robust zero-shot VC.
@@ 3.3.1 Attribute-Conditional Normalizing Flow for Age and Gender Editing
We observe that the speaker encoder jointly trained with our diffusion backbone model achieves strong speech attribute disentanglement, indicating that many attributes like age and gender are encoded into the global 512-dimensional speaker embedding vector. Therefore, the manipulation of these attributes can be seen as re-sampling from the learned latent space of speaker embeddings. To achieve fully controllable generation and editing of specific speaker attributes while leaving other attributes unaffected, we take inspiration from StyleFlow (Abdal et al., 2021) in the image editing domain and employ a similar attribute-conditional continuous normalizing flow (CNF) that operates on this latent space.
@@
The training objective of the CNF would be to maximize the likelihood of the data $\mathbf{w}$. In our attribute conditional scenario, given the speaker attributes vector $\mathbf{a}$ (e.g., for age and gender, $\mathbf{a}\in\mathbb{R}^{2}$) that is associated with sample $\mathbf{w}$, we update the ODE network to be $\phi_{\theta}(z(t),\mathbf{a},t)$, which is conditioned on both $t$ and $\mathbf{a}$. Consequently, our new objective becomes $\max_{\theta}\sum_{\mathbf{w},\mathbf{a}}\log p(\mathbf{w}|\mathbf{a},t)$. During training, we also employ trajectory polynomial regularization (Huang and Yeh, 2021), which we find stabilizes the training. Note that there is no attribute editing during the training process. A well-trained CNF model behaves as a bijective mapping between the data distribution and prior distribution with the help of an ODE solver. In our experiments, we use the dopri5 ODE solver (Hairer et al., 2008) and adopt the same CNF implementation from StyleFlow [^3] https://github.com/RameenAbdal/StyleFlow. We only use one CNF block with hidden dimension 512 rather than stacking multiple blocks, which leads to lower validation loss. Our lightweight CNF model has 0.5M learnable parameters and is trained using Adam (Kingma and Ba, 2014) on a single V100 GPU for 24H. The attribute editing procedure at inference time is depicted in Figure 2.
Attributes Dataset.
The attribute labels for training our attribute-conditional CNF are obtained from the CommonVoice 13.0 dataset (Ardila et al., 2019), which contains multiple audio recordings of more than 51K human-validated anonymous speakers, totaling 2,429 hours. Among these validated speakers, about 20K speakers provide age (from twenties to nineties) and gender labels (male or female). To fully utilize the dataset, we first train a naive ECAPA-TDNN-based age and gender prediction model by appending a projection layer that predicts age as a numeric value and gender as a logit to its original output layer. We combine mean absolute error (MAE) and cross-entropy (CE) losses for age and gender prediction respectively. This prediction model serves two purposes: